OpenStudy (anonymous):

How do I find the limit of x/x-4 as x->4 without plugging in numbers close to 4?

OpenStudy (anonymous):

If its \[\lim_{x\rightarrow 4}{\frac{x}{x-4}}\] Then the answer is undefined! as in the answer is standard for a reciprocal question: \[\lim_{x\rightarrow 4^+}{\frac{x}{x-4}}=\infty\phantom{pad to next question}\lim_{x\rightarrow 4^-}{\frac{x}{x-4}}=-\infty\] But if it's: \[\lim_{x\rightarrow4}{\left[\frac{x}{x}-4\right]}\] Then in general: \[\lim_{x\rightarrow a}{\left[\frac{x}{x}-4\right]}=-3\]

OpenStudy (anonymous):

It's the first one. How do I show that it is infinite from the right or from the left without plugging in numbers?

OpenStudy (anonymous):

Erm...without numbers...hmm

OpenStudy (anonymous):

Im not quite sure to be honest, I wouldn't want to lie haha

OpenStudy (anonymous):

Alright, thanks anyway!